OpenStudy (anonymous):

Is the line through (4,1,-1) and (2,5,3) perpendicular to the line through (-3,2,0) and (5,2,4)? please justify

OpenStudy (anonymous):

Suppose a line through point A and B, a second line through C and D: U = B-A V = D-C if U dot V = 0 then perpendicular, as X dot Y = |X|*|Y|*Cos(t) where t is 1angle between the two vectors